### v1

# CART: Caltech Aerial RGB-Thermal Dataset in the Wild

Abstract
We present the first publicly available RGB-thermal dataset designed for aerial robotics operating in natural environments. Our data-set captures a variety of terrains across the continental United States, including rivers, lakes, coastlines, deserts, and forests, and consists of synchronized RGB, long-wave thermal, global positioning, and inertial data. Furthermore, we provide semantic segmentation annotations for 10 classes commonly encountered in natural settings in order to facilitate the development of perception algorithms robust to adverse weather and nighttime conditions. Using this dataset, we propose new and challenging benchmarks for thermal and RGB-thermal semantic segmentation, RGB-to-thermal image translation, and visual-inertial odometry. We present extensive results using state-of-the-art methods and highlight the challenges posed by temporal and geographical domain shifts in our data. Dataset and accompanying code will be provided at https://github.com/aerorobotics/caltech-aerial-rgbt-dataset.
Keywords:
Robotics Semantic segmentation RGB-T image translation Visual inertial odometry
**footnotetext: These authors contributed equally to this work
Figure 1: Left: Our dataset is the first dataset designed to improve thermal semantic perception for field robotics. Right: Our dataset provides new thermal and RGB-thermal benchmarks for field robotics and computer vision algorithms, encompassing (a) semantic segmentation, (b) image translation, and (c) motion tracking.
1 Introduction
Current field robots rely predominantly on sensors such as visual cameras, lidar, and radar to perceive their surroundings [16, 18]. While these sensors enhance robustness of downstream vision algorithms, their performance degrades in low-light and adverse weather (snow, fog, rain) conditions [68]. In contrast, thermal cameras exploit long-wave infrared wavelengths to capture emitted heat, offering dense, visual information even in conditions in which other methods struggle [19]. Recently, thermal cameras have been used in multimodal perception algorithms for autonomous vehicle applications and are increasingly being explored for field robotic applications to enable nighttime autonomy [36, 74, 69, 11, 76, 48, 12]. However, successful integration of thermal imagery requires extensive datasets across diverse settings, preventing widespread adoption in field robotics.
Existing thermal datasets primarily focus on urban environments for autonomous driving applications (Fig. 1). They typically comprise of thermal-only imagery [74, 36] or RGB-Thermal (RGB-T) pairs [17, 22, 69, 40], with some including global positioning (GPS/GNSS) and inertial measurements (IMU) for visual-inertial odometry (VIO) and simultaneous localization and mapping (SLAM) [76, 11, 60]. While comprehensive, these datasets rarely extend beyond urban areas. They lack data depicting natural settings like rivers and forests, which are typical operating areas for field robotics that perform coastline mapping and bathymetry [4] or monitor activity during forest fires [30].
Due to the lack of thermal benchmarks for algorithms like semantic segmentation and SLAM, field robots cannot easily operate at night or in adverse conditions, requiring online learning [70, 34] or unsupervised domain adaptation to compensate [20, 32, 69] which still require labeled data for evaluation. As such, enabling thermal perception for field robotics requires collecting and annotating field-specific data from scratch. However, curating such a dataset is more complex than doing so for RGB: First, thermal data is hard to crowdsource and web-scrape due to the cost and nicheness of thermal cameras. Hence, acquiring a dataset for algorithm development requires physically going to distinct field environments, increasing time and cost. Second, operating most field robotics, especially uninhabited aerial vehicles (UAV), usually require special permits unique to each location-of-capture. This further complicates the creation of a comprehensive thermal dataset for field robotics research.
In this work, we present the first dataset targeting thermal scene perception for field robotics and for wider use by the computer vision community. The dataset consists of oblique-facing imagery captured from a UAV, supplemented with image sets captured at ground level, and focuses on littoral settings within various desert, forest, and coastal environments across the United States. We contribute new benchmarks for thermal and RGB-T semantic segmentation, RGB-T image translation, and VIO/SLAM algorithms, with unique challenges characterized by temporal and geographical domain shift for learning-based methods and periodic feature sparsity for motion tracking algorithms.
The paper is organized as follows: Sec. 2 reviews relevant datasets, benchmarks, and algorithms. Sec. 3 describes our dataset details and curation process, and Sec. 4 presents benchmark results. Sec. 5 offers concluding remarks.
2 Related Work
2.1 Datasets and Benchmarks
2.1.1 Thermal/RGB-T benchmarks for urban robotics
Current urban thermal and RGB-T datasets primarily focus on autonomous vehicle (AV) technology and surveillance applications. AV-related datasets cover tasks such as object detection [17, 40, 11], semantic segmentation [36, 22, 69, 74], and localization [76, 11, 60]. Some include data from other spectra, benchmarking not only thermal algorithms but also RGB-T [40, 11, 22, 69] and other multispectral models. In contrast, datasets for surveillance applications use fixed cameras [29] and UAVs [65] to detect personnel, vehicles, and other objects. Despite many urban thermal benchmarks, they cannot be directly used to develop algorithms for robots operating in natural environments due to the urban/non-urban domain gap.
2.1.2 Thermal/RGB-T benchmarks for field robotics
Datasets for robots operating in non-urban environments remain limited. MassMIND [48] and PST900 [61] benchmark thermal semantic segmentation for maritime and subterranean robotic environments, respectively. WIT-UAS [30] provides UAV-borne thermal images for object detection in wildfire-prone environments. [34] proposes an aerial RGB-T dataset for thermal water segmentation in littoral areas but only provides 1272 labels consisting of a single water class. Benchmarks like [3, 7] depict natural environments but are targeted towards wildlife tracking. To address this gap, we present an aerial RGB-T dataset which can be used to benchmark multiple tasks, including semantic segmentation, RGB-T image translation, and localization, across diverse natural environment robotic applications. Specifically, we build upon the dataset from [34] by extending the number of semantic classes from 1 to 10 and adding additional RGB-T data captured from air and ground, resulting in a total of 4195 semantic segmentation annotations.
2.1.3 UAV-based semantic segmentation datasets
Aside from [34], existing UAV-captured semantic segmentation datasets [47, 10, 44, 6, 46] cover urban environments. Wth the exception of [46] which provides 400 thermal image samples, all of these datasets contain only RGB data. Although they provide images with similar view angles as in our dataset, the RGB-T modality gap and their focus on urban semantic content makes them impractical for field robotic applications.
2.2 Algorithms using Thermal Datasets
2.2.1 Thermal semantic segmentation
Since thermal images are spatially identical to their RGB counterparts, popular RGB semantic segmentation algorithms [52, 8, 5, 72] can be applied directly. However, some works [36, 74, 50] develop models specifically for thermal imagery, and leverage edge priors to overcome low resolution and blurriness due to thermal crossover [55]. Although specifically designed for thermal, these models only marginally outperform standard RGB segmentation models [50] and their performance in field settings, where edges are less structured, remains untested. Aside from architectural advancements, improvements in thermal semantic segmentation come from domain adaptation (DA) methods [20, 36, 69, 67] that use labeled RGB data with unlabeled thermal data. To provide fair comparisons, we do not provide DA baselines in our benchmarks as performance varies based on the choice of RGB data.
2.2.2 RGB-T semantic segmentation
RGB-T semantic segmentation methods compensate for the weaknesses of each modality via deep feature fusion, data augmentation, and adversarial training. Most approaches [22, 64, 63, 59, 77] utilize encoder-decoder architectures with modality-specific encoders and shared decoders, and show significant improvements over channel-stacked RGB-T inputs. Works like [38, 79] make improvements by integrating RGB and thermal features at multiple stages within the encoders, while [61] eschews RGB-T training pairs by processing RGB-T inputs sequentially. Other methods include random input masking to reduce reliance on a single modality [59] and adversarial training to adapt to different times of day [69].
2.2.3 RGB-T image translation
RGB-T image translation is vital as it may enable thermal training data to be generated from large-scale RGB datasets. Works based on generative adversarial networks (GAN) such as [41, 27, 80, 51, 35] can be used without RGB-T image pairs. However, GANs like [28, 71] that require paired imagery provide better translations for both visual appeal and domain adaptation [36]. Recent diffusion methods [57, 56] have demonstrated remarkable results on RGB imagery, but have not been tested on RGB-T image translation.
2.2.4 Thermal VIO and SLAM
Low signal-to-noise ratio, reduced contrast, and blurred boundaries in thermal images present challenges for VIO and SLAM. Several studies [2, 13, 24, 26, 14] use indirect methods by tracking keypoints extracted from thermal images for visual constraints and rely on image normalization techniques to improve keypoint extraction. In contrast, [31] uses a direct method to extract and track high-gradient points by minimizing radiometric error. Other works utilize learned local and global features to improve feature tracking [78] and loop closure detection [58], respectively.
3 The Aerial RGB-Thermal Dataset
3.1 Data Acquisition Hardware
[…]
Figure 2: (a) The Aurelia X6 hexacopter and sensor stack used to capture our aerial dataset. (b) Geographic distribution of our data collection sites and collection times.
[…]
3.2 Data Collection and Processing
3.2.1 Data capture
[…]
3.2.2 Thermal image normalization
For the baselines presented in this work, we typically normalize 16-bit thermal data by rescaling between the 1st and 99th percentile pixel values and follow with contrast limited adaptive histogram equalization (CLAHE). We keep normalized data as floats to minimize information loss, and only convert to 8-bit format when necessary or for visualization.
Figure 3: Semantic segmentation classes in our dataset. The color mapping is used throughout this paper. (a) Hourly distribution of annotated thermal images. (b) Histogram of semantic classes.
3.2.3 Semantic segmentation annotation
[…]
3.2.4 RGB-thermal image alignment
To align thermal and RGB image pairs, we stereo rectified the RGB-T image pairs using the camera matrices obtained from calibration (Sec. S2.2) and projected the thermal image into the larger RGB image frame to preserve RGB resolution. As our baseline is small compared to the depth of the scenes we capture, image pairs are near-coregistered after rectification. Due to differences in optics, the field-of-view of the aligned image pairs is narrower compared to that of the thermal camera.
3.3 Dataset Splits
3.3.1 General (benchmark) split
This is our primary split for semantic segmentation and image translation benchmarks. We randomly partition the annotated thermal dataset (Sec. 3.2.3) into train/val/test sets at a 75:12.5:12.5 ratio.
3.3.2 Temporal split
To promote studies into the effect of different time-of-day capture of the thermal images, we split the dataset using three time periods: twilight/sunrise (5 AM - 7 AM), daytime (10 AM - 5 PM), and nighttime (7 PM - 4 AM). We randomly partition the daytime images into train/val/test sets using the ratio from above and leave the twilight and nighttime sets for testing.
3.3.3 Geographical split
The geographical splits are intended to test algorithm adaptability to unseen settings. We provide two splits:
  1. 1.
    Terrain-based: This split partitions data into the following terrain categories: aerial river, aerial lake, aerial coast, and ground-captured images. Train/val/test splits are created per category with a 50:15:35 ratio.
  2. 2.
    Region-based: This split partitions data based on the general area of the United States that the data was captured. The regions include California (CA), Kentucky (KY), and North Carolina (NC), with train/val/test splits created in the same manner as the terrain split.
4 Experiments
We provide baselines for the following benchmarks: thermal semantic segmentation, RGB-T semantic segmentation, RGB-T image translation, and VIO/SLAM. Furthermore, we explore the zero-shot thermal segmentation capabilities of large vision-language models on our dataset and perform further ablation studies.
Table 1: Supervised model performance (mIoU), framerate (FPS at batch size 1), and floating point operations (FLOP) for various semantic segmentation network baselines.
Model
Bare ground
Rocky terrain
Develop. struct.
Road
Shrubs
Trees
Sky
Water
Vehicles
Person
Stuff
Object
All
Orin (GPU)
NUC (CPU)
FLOPS (G)
FastSCNN [52]
0.807
0.907
0.728
0.631
0.694
0.776
0.956
0.965
0.384
0.058
0.808
0.221
0.690
124.2 85.6 1.0
MobileNetV3-S 0.75 [25]
0.802
0.914
0.676
0.633
0.704
0.814
0.955
0.976
0.383
0.111
0.809
0.247
0.697
75.2
24.0
3.3
MobileNetV3-L 0.75
0.801
0.914
0.701
0.615
0.713
0.793
0.959
0.976
0.432
0.148
0.809
0.290
0.705
59.0
15.5
4.8
MobileViTV2 0.50 [45]
0.822
0.913
0.759
0.650
0.694
0.781
0.954
0.970
0.222
0.030
0.818
0.126
0.680
60.1
10.7
5.5
EfficientViT-B0 [5]
0.825
0.917
0.777
0.618
0.710
0.793
0.961
0.965
0.495
0.191
0.821
0.343 0.725
51.2
5.5
3.9
EfficientNet-Lite0 [66]
0.819
0.908
0.769
0.601
0.709
0.805
0.962
0.968
0.380
0.079
0.818
0.229
0.700
47.7
12.8
7.2
EfficientNet-Lite2
0.825
0.920
0.765
0.614
0.698
0.805
0.958
0.971
0.386
0.118
0.820
0.252
0.706
38.9
11.2
9.4
Segformer-B0 [72]
0.804
0.910
0.690
0.624
0.696
0.788
0.960
0.970
0.195
0.000
0.805
0.097
0.664
38.3
5.9
10.2
Segformer-B1
0.814
0.909
0.773
0.603
0.713
0.799
0.960
0.965
0.366
0.000
0.817
0.183
0.690
30.3
3.5
19.9
ResNet18 [23]
0.810
0.905
0.766
0.618
0.711
0.807
0.959
0.966
0.431
0.158
0.818
0.294
0.713
69.3
12.4
22.4
ResNet50
0.819
0.916
0.747
0.658
0.711
0.799
0.961
0.974
0.361
0.182
0.823
0.272
0.713
29.9
6.1
45.4
ResNeXt50 [73]
0.825
0.919
0.773
0.653
0.709
0.794
0.964
0.976
0.436
0.137
0.827
0.287
0.719
23.0
5.6
45.5
ConvNext-T [43]
0.799
0.909
0.719
0.608
0.706
0.808
0.961
0.970
0.363
0.003
0.810
0.183
0.685
25.8
3.3
47.4
ConvNext-S
0.810
0.913
0.771
0.603
0.718
0.810
0.965
0.964
0.492
0.048
0.819
0.270
0.709
19.1
2.3
75.0
ConvNext-B
0.810
0.921
0.697
0.611
0.706
0.803
0.963
0.972
0.348
0.112
0.811
0.230
0.694
14.6
1.4
130.5
ConvNext-B (CLIP) [54]
0.813
0.918
0.683
0.632
0.718
0.813
0.965
0.972
0.517
0.137
0.814
0.327
0.717
14.6
1.4
130.4
ConvNext-B (CLIP) ❄
0.773
0.887
0.713
0.487
0.656
0.756
0.946
0.951
0.158
0.019
0.771
0.089
0.635
14.6
1.4
130.4
DINOv2 [49] (linear head)
0.800
0.907
0.681
0.606
0.693
0.796
0.956
0.967
0.375
0.079
0.801
0.227
0.686
15.2
2.1
DINOv2 (linear head) ❄
0.705
0.844
0.619
0.419
0.558
0.725
0.922
0.920
0.356
0.158
0.714
0.257
0.623
15.2
2.1
DINOv2 (nonlin. head)
0.810
0.916
0.708
0.635
0.700
0.819
0.959
0.973
0.399
0.119
0.815
0.259
0.704
15.2
2.1
DINOv2 (nonlin. head) ❄
0.796
0.903
0.732
0.606
0.691
0.794
0.954
0.963
0.471
0.149
0.805
0.310
0.706
15.2
2.1
${}^{\dagger}$ [50]
0.755
0.867
0.635
0.576
0.643
0.653
0.787
0.947
0.234
0.024
0.733
0.129
0.613
11.1
1.4
100.5
${}^{\dagger}$hermal-specific model  ❄ Frozen encoder
4.1 Thermal Semantic Segmentation
We run multiple baselines on our general split before choosing a specific model to test on the geographic and temporal splits. We go over our baselines before analyzing their performance on our benchmarks.
4.1.1 Baselines
Motivated by robotic applications, we are interested in both inference speed and segmentation performance. Our baselines place an emphasis on lightweight, real-time networks, but also on foundation models like DINOv2 [49] and ConvNext-B (CLIP) [43, 54] to explore any latent multimodal capabilities. We also test a thermal-specific model, FTNet [50], which uses an edge-based loss function to compensate for blurred boundaries in thermal images.
Besides FastSCNN [52], EfficientViT [5], Segformer [72], DINOv2, and FTNet, all network encoders in Tab. 1 feed into a DeepLabV3+ segmentation head. In particular, DINOv2 was employed with linear and non-linear, multi-scale heads in order to study the generalization capacity of its pretrained features to the thermal modality. All baselines except DINOv2 and ConvNext-B (CLIP) were trained from pretrained ImageNet weights. All networks, besides FTNet, were trained using the cross entropy loss. Further details on baseline implementation and training can be found in Supplementary Material Sec. S3.1.
Figure 4: Thermal images and semantic segmentation labels from each capture area with inference results from EfficientViT, FastSCNN, and ConvNext-B (CLIP).
4.1.2 General benchmark
We use the general (random) split (Sec. 3.3.1) to compare thermal semantic segmentation performance between baselines (Tab. 1). With field robotic applications in mind, we analyze segmentation results in context of GPU and CPU frame rates, which were measured onboard the Nvidia Jetson AGX Orin (GPU) and the Ruby NUC (CPU) embedded platforms, respectively [^2] Models of input size 512$\times$640 were converted to ONNX and inference speed was averaged over 50 forward passes after a warm-up period..
[…]
While large RGB foundation models show little gain over ImageNet pretraining, their features do extend to the thermal modality. Specifically, a frozen DINOv2 (linear head) outperforms the thermal-specific FT-Net. With a nonlinear head, it matches the performance of other end-to-end trained networks. While this shows the benefit of large-scale RGB pretraining, the choice of RGB pretraining data or learning strategy seems to matter. Notably, the frozen ConvNext-B (CLIP) model gives marginal gain over the frozen DINOv2 (linear head), despite a larger, nonlinear segmentation head. Overall, foundation models do not confer significant performance advantages for semantic segmentation on our dataset in order to justify their computational cost.
Overall, all baselines can segment stuff classes well but none excel with the rare object classes: vehicle and person. Future models would likely need to consider weighted loss functions, smarter data augmentation, or domain adaptation techniques in order to improve on object classes. Lastly, we note that the thermal-specific model, FTNet, performs poorly, likely as its edge priors are less effective in unstructured, natural environments compared to urban environments.
Figure 5: (a) Thermal semantic segmentation failures due to intra-class semantic variations when testing on geographically-partitioned, out-of-domain data. (b) Failures due to photometric variations between day (in-domain) and night (out-of-domain).
4.1.3 Geographically-partitioned benchmarks
We retrained our best baseline
(Tab. 1) on geographically-partitioned sets to evaluate intra-thermal generalization. Given the cost and effort of collecting diverse aerial thermal data of field settings, we aim to determine the least amount of annotated data required to effectively deploy. We performed two experiments using EfficientViT-B0: one on generalization across geographic regions and another across differing terrain types.
Table 2: Thermal semantic segmentation results on geographically-split data
(a) Region-based split
Train Area Test Area Avg. mIoU
CA
NC
KY
CA
0.666
0.084
0.193
0.314
NC
0.084
0.508
0.031
0.208
All
0.648
0.525
0.587
0.586
(b) Terrain-based split
Training Terrain Testing Terrain Avg. mIoU
River
Lake
Coast
River
0.668
0.196
0.117
0.327
Lake
0.196
0.364
0.061
0.207
Coast
0.127
0.119
0.522
0.256
All
0.646
0.415
0.493
0.518
All + ground
0.658
0.416
0.520
0.532
In the first experiment, we trained on thermal data from CA and tested on images from KY and NC (Tab. 2(a)). Despite common classes in each region, results show poor generalization to class variations across geographic areas. We see this again when repeating with the NC split. In the second experiment, we divided our data into three terrain types, aerial river, aerial lake, and aerial coast. Remaining ground-captured images were lumped into a fourth ground category. We trained and tested all pairwise combinations and found poor generalization across terrain types as well (Tab. 2(b)). Notably, we found low, in-domain performance for the lake setting but saw gains after training on aerial data from other terrain and further improvements after training on ground level images. Overall, our results reiterate the benefit of having training data representative of the testing environment, but also show that gains can be made by including out-of-domain (OOD) data from different terrain and different view points.
Overall, the geographic benchmark presents a difficult domain adaptation problem, requiring both inter- and intra-modality domain adaptation techniques to overcome. Although such techniques could perform well in this challenge, the results also highlight a need for an even larger and more diverse thermal dataset capturing common field environments than what we presented.
4.1.4 Temporally-partitioned benchmark
Here, we assess EfficientViT-B0’s performance against intra-thermal, temporal distribution shift due to thermal inversion. We train the model on images from the daytime and test on images captured at sunrise and at night. The model achieved mIoUs of 0.242 at sunrise, 0.193 at night, and 0.777 on a separate daytime test split. Despite training on daytime images from the same location, the model fails on nighttime scenes, struggling to classify water due to thermal inversion (Fig. 5b). As such, to perform well on this split, future algorithms need to augment datasets to simulate such inversions or intentional collect data capturing such phenomena.
Table 3: RGB-Thermal semantic segmentation network baseline results (mIoU).
Model
Bare ground
Rocky terrain
Devel. struct.
Road
Shrubs
Trees
Sky
Water
Vehicles
Person
All
FLOPS (G)
RGB ${}^{\dagger}$
0.790
0.873
0.808
0.559
0.664
0.752
0.891
0.967
0.399
0.000
0.670
45
Thermal ${}^{\dagger}$
0.766
0.835
0.789
0.536
0.654
0.749
0.907
0.971
0.399
0.000
0.661
45
EAEFNet [38]
0.805
0.863
0.813
0.545
0.702
0.802
0.937
0.976
0.481
0.000
0.692
358
CRM [59]
0.816
0.900
0.851
0.616
0.695
0.779
0.923
0.974
0.377
0.206
0.714
310
CMNeXt [77]
0.830 0.900 0.861
0.607
0.718 0.808
0.933
0.980 0.560
0.190
0.740
148
${}^{\dagger}$ DeepLabV3+ (w/ ResNet50 encoder)
4.2 RGB-T Semantic Segmentation
We use the paired RGB-T dataset (Sec. 3.2.4) and partition it using the general split (Sec. 4.1.2). Recall that RGB and thermal input sizes are 960$\times$600 pixels and the thermal FoV is narrower than in the previous benchmarks (Sec. 4.1).
4.2.1 Baselines
We test three RGB-T semantic segmentation algorithms achieving state-of-the-art results in urban RGB-T scene segmentation benchmarks: CRM [59], EAEFNet [38], and CMNeXt [77]. To compare against single-modality models, we train two additional ResNet50/DeepLabV3+ segmentation models on RGB and thermal data accordingly (Sec. 4.1).
4.2.2 Performance analysis
All RGB-T models outperform the single-modality baselines, with CMNeXt outperforming the next best RGB-T model by 0.026 mIoU (Tab. 3). Overall, incorporating RGB imagery into the segmentation pipeline greatly improved the distinction between land-based stuff classes which can be hard to distinguish in thermal due to thermal crossover. However, these improvements come at a high computational cost and would not currently be suitable for field robotic applications. Excelling in this benchmark would involve improving segmentation performance of the object classes, possibly by leveraging RGB data in unsupervised domain adaptation training schemes, as well as increasing computational efficiency through prudent architectural design choices.
4.3 RGB-Thermal Image Translation
[…]
4.3.1 Baselines
We benchmark GAN methods that need RGB-T pairs (Pix2Pix [28], Pix2PixHD [71], VQGAN [15]) and unpaired methods (UNIT [41], MUNIT [27], [35]) that do not. We also evaluate Palette [57], a paired diffusion approach.
Figure 6: RGB-to-thermal image translation results. Zoom-in to see fine-details.
4.3.2 Image translation metrics
We quantify RGB-T image translation using the Peak-Signal-to-Noise ratio (PSNR) and the Structural Similarity Index (SSIM). In addition, we propose a third metric: the thermal mIoU. This is equivalent to the FCN-score used in RGB image translation works [28]. Instead of an RGB network, our metric uses our EfficientViT thermal segmentation network (Tab. 1) to segment translated thermal images before evaluating with ground truth.
Table 4: RGB-Thermal image translation results
Method Method Type RGB $\rightarrow$ Thermal
PSNR $\uparrow$
SSIM $\uparrow$
mIoU $\uparrow$
UNIT [41]
Unpaired GAN
12.85
0.421
0.189
MUNIT [27]
Unpaired GAN
16.08
0.459
0.173
Edge-guided RGB-T [35]
Unpaired GAN
12.20
0.436
0.225
Pix2Pix [28]
Paired GAN
20.06
0.547
0.354
Pix2PixHD [71]
Paired GAN
20.05
0.579
0.379
VQGAN [15]
Paired GAN
18.09
0.540
0.388
Palette [57]
Paired Diffusion
11.51
0.399
0.194
4.3.3 Performance analysis
Paired GANs outperform unpaired GANs and diffusion methods, achieving higher values across all three metrics (Tab. 4). Qualitative assessment (Fig. 6) shows poor translations from unpaired techniques, with most retaining geometric artifacts unique to RGB (ocean waves and shadows) and ignoring relative temperature characteristics. Paired GANs produce results that appear similar to real thermal images but with inconsistent details upon closer inspection. Likewise, the diffusion method produces accurate translations but is not consistent. Overall, RGB-T translation still requires improvements before being able to provide a reliable source of thermal training data from existing RGB datasets. For domain adaptation purposes, we emphasize that maximizing the thermal mIoU score is more important than photometric consistency.
4.4 Motion Tracking
To quantify VIO/SLAM robustness, we select 12 clipped sequences from our dataset ranging in motion tracking difficulty: from urban sports fields (easy) to our natural environments (hard). We normalize thermal images using the 5th and 95th percentile pixel values (Sec. 3.2.2) to enable feature detection.
Table 5: VIO/SLAM performance (Absolute Trajectory Error [m]) on aerial sequences
Method Type Modality North Field Castaic Lake, CA Duck, NC
N1
N2
N3
N4
N5
N6
C1
C2
C3
C4
D1
D2
VINS-Fusion
SLAM
RGB
2.883
7.530
4.036
2.009
4.045
0.834
1.555
3.277
1.566
1.031
0.700
VINS-Fusion
Thermal
9.454
12.26
9.854
2.926
11.00
5.114
7.296
1.513
5.321
2.025
0.879
VINS-Fusion
VIO
RGB
5.458
9.509
6.518
2.121
6.026
1.277
1.555
3.277
1.377
1.031
0.725
VINS-Fusion
Thermal
13.82
12.86
11.99
2.926
13.98
5.194
7.284
10.60
5.321
2.034
0.879
Open-VINS
VIO
RGB
22.80
36.15
37.15
5.165
3.896
3.573
0.475
0.700
Open-VINS
Thermal
14.43
30.64
16.12
2.258
1.562
5.311
1.073
Trajectory length (m)
1533
1326
1253
153
1104
642
310
65
174
137
88
80
4.4.1 Baselines
We evaluate VINS-Fusion [53] (graph optimization-based) and Open-VINS [21] (filtering-based). We test VINS-Fusion in VIO and SLAM modes, with and without loop closure constraints, respectively, and report Absolute Trajectory Error [62] averaged over four runs.
4.4.2 Performance analysis
This benchmark reveals challenges due to fast motion at altitude and periodic feature sparsity in littoral environments. In such scenes, motion tracking frequently failed due to loss of features, requiring sequences to be clipped for quantifiable evaluations (Tab. 5). VINS-Fusion outperformed Open-VINS mainly due to higher feature tracking reliability. VINS-Fusion works better on RGB images than on thermal images in general. However, VINS-Fusion demonstrates superior feature tracking on thermal images compared to RGB on the C2 sequence of Castaic Lake, where the thermal contrast of stone grains on the lake coast was particularly pronounced in the late afternoon. This observation motivates future research on integrating thermal and RGB cameras for motion tracking with enhanced robustness and versatility. Our benchmark presents a difficult challenge for VIO and SLAM algorithms in extreme environments where water and reflections off water surfaces dominate image scenes, providing a unique test bed to invigorate interest in robust feature matching and motion tracking in natural texture-deficient scenarios.
4.5 Further Analysis
4.5.1 Zero-shot foundation models on thermal imagery
Current foundation models work well on RGB imagery but have not been tested in the thermal domain, especially in non-urban environments. To evaluate foundation models on thermal imagery, specifically for zero-shot segmentation and semantic segmentation, we compare RGB and thermal performance using the aligned set (Sec. 3.2.4). With this set, we isolate modality as the root cause for any performance difference. In the following experiments, our metrics do not include the rare object classes (vehicle and person) as they drop metrics to near-uniform values for both modalities, making it difficult to compare.
Figure 7: Zero-shot foundation models on RGB-T pairs. Oversegmented SAM outputs differ across modalities but are similar with ground truth semantics added. Semantic segmentation models (prompted with all classes) perform better on RGB imagery.
Table 6: RGB foundation model performances on thermal imagery
(a) Instance segmentation via SAM on thermal and augmented color images using coregistered color image segmentations as ground truth
Modality
[…]
Thermal
0.018
0.007
0.022
0.204
Color (grayscale)
0.538
0.536
0.531
0.601
Color (col. jitter)
0.800
0.792
0.801
0.841
${}^{\dagger}$refers to AP@0.5::0.95, following MSCOCO [39]
(b) Zero-shot semantic segmentation (mIoU) on coregistered color and thermal imagery
Method Color
Color (gray)
Color (jitter)
Thermal
SAM [33] + GT ${}^{\ddagger}$
0.704
0.698
0.700
0.653
Grounded SAM [42]
0.380
0.351
0.361
0.193
OV-Seg [37]
0.362
0.340
0.365
0.217
ODISE [75]
0.504
0.450
0.486
0.232
${}^{\ddagger}$ Most frequent ground truth class label in a SAM mask assigned as label for entire mask
We quantify Segment Anything’s (SAM) [33] zero-shot segmentation sensitivity to modality change by measuring the spatial alignment of its mask outputs. We apply it to the RGB set and the 8-bit thermal image set (Sec. 3.2.2) with the default 32$\times$32 grid points to segment everything (Fig. 7). We compute the average precision (AP) of the predicted thermal masks, using the predicted RGB masks as ground truth, and compare the thermal AP against the APs of a color-jittered RGB set and a grayscale set. Results show that SAM is sensitive to photometric changes, with sizable performance drops from color-jitter to grayscale to thermal (Tab. 6(a)).
Next, we test large vision-language models (Grounded-SAM [42], OV-Seg [37], and ODISE [75]) in zero-shot semantic segmentation. We prompt the models with a list of classes in our dataset; generate semantic segmentation masks for RGB, color-jittered RGB, grayscale, and thermal image sets; and compute the mIoU of their outputs using our ground truth annotations. Although RGB, RGB-variants, and thermal modalities all yield poor semantic segmentation performance (Tab. 6(b)), all RGB variants outperform thermal by at least 0.13 mIoU, indicating low generalization to the thermal modality (Fig. 7).
Finally, we create semantic SAM masks by assigning class labels to SAM instances based on the most frequent ground truth class within an instance (Fig. 7). We find that semantic SAM masks attain much higher mIoU score (0.653) compared to the other zero-shot methods. Based on the overall results (Tab. 6), we conclude that SAM can delineate object boundaries in the thermal domain reasonably well but can oversegment instances resulting in poor AP scores when compared to RGB (which can also oversegment). When paired with ground truth semantic labels, this behavior is hidden.
4.5.2 Transfer learning with urban thermal datasets
Table 7: Network pretraining methods for downstream thermal segmentation.
[…]
mIoU # Params. (M)
None
ImageNet
Therm. Urban
[…]
0.640
0.690
0.688
1.1
EfficientViT
0.687
0.725
0.714
4.8
ResNet18
0.702
0.713
0.706
12.3
ResNet50
0.682
0.713
0.728
26.7
ConvNext-B
0.625
0.717
0.697
89.4
[…]
5 Conclusion
We presented the Aerial RGB-Thermal Dataset, the first publicly available dataset specifically tailored towards advancing thermal semantic perception and motion tracking algorithms in natural environments. We created four distinct benchmarks encompassing semantic segmentation in both thermal and RGB-T domains, RGB-T image translation, and motion tracking, and use these benchmarks to demonstrate the current challenges faced by thermal-based robot perception and localization algorithms. Current semantic segmentation and image translation methods were particularly affected by geographical domain shifts, reflecting the diverse intra-class and inter-scenery distributions in our dataset, as well as temporal domain shifts due to thermal inversion and crossover. Motion tracking algorithms, on the other hand, suffered from poor feature tracking in the thermal modality due to reduced spatial quality as compared to RGB and encountered failures when faced with challenging, feature-sparse scenarios. However, such scenarios are commonly encountered during real-world deployments with failures of significant cost and consequence [1]. Our motion tracking benchmark greatly penalizes current VIO/SLAM algorithms for assuming ideal feature-tracking conditions, serving as a unique testbed for algorithm improvements in this area. This dataset, along with the benchmarks, can be used into the future to help develop algorithms that help expand the operating domains of both robotics and computer vision in general.